### v1

# Online Optimization for Learning to Communicate over Time-Correlated Channels

Abstract
Machine learning techniques have garnered great interest in designing communication systems owing to their capacity in tacking with channel uncertainty. To provide theoretical guarantees for learning-based communication systems, some recent works analyze generalization bounds for devised methods based on the assumption of Independently and Identically Distributed (I.I.D.) channels, a condition rarely met in practical scenarios. In this paper, we drop the I.I.D. channel assumption and study an online optimization problem of learning to communicate over time-correlated channels. To address this issue, we further focus on two specific tasks: optimizing channel decoders for time-correlated fading channels and selecting optimal codebooks for time-correlated additive noise channels. For utilizing temporal dependence of considered channels to better learn communication systems, we develop two online optimization algorithms based on the optimistic online mirror descent framework. Furthermore, we provide theoretical guarantees for proposed algorithms via deriving sub-linear regret bound on the expected error probability of learned systems. Extensive simulation experiments have been conducted to validate that our presented approaches can leverage the channel correlation to achieve a lower average symbol error rate compared to baseline methods, consistent with our theoretical findings.
Index Terms:
Time-correlated channels, decoder learning, codebook selection, online optimization theory, online convex optimization, multi-armed bandit, error probability analysis.
I Introduction
The widespread adoption of machine learning techniques in developing communication systems based on real-world data has sparked broad interest in recent years [1]. Machine learning algorithms have shown to be effective tools for various tasks such as channel estimation [2], equalization [3], coding [4], decoding [5], and other physical layer applications [1, 6]. Learning-based communication systems have showed impressive performance in their capacity to generalize effectively to unknown channels [1, 7].
Most existing studies utilizing machine learning approaches to design communication systems lack theoretical justification for their proposed methods, and typically regard learned communication systems as black boxes [6, 7, 8]. Recently, a few works are aiming to conduct theoretical analyses for learning-based communication systems [9, 10, 11]. Specifically, these studies leverage statistical learning theory [12, 13] to derive generalization bounds based on the Independently and Identically Distributed (I.I.D.) channel assumption. Building upon this assumption, the generalization capacity of learning-based communication systems can be assured. This is because these systems can be trained using sufficient data during the offline stage and demonstrate commendable performance when evaluated with data sampled from the same distribution.
Figure 1: The procedure of learning decoder and codebook using online optimization methods.
However, meeting the assumption of I.I.D. channels in reality is generally challenging, given that practical communication scenarios often involve time-varying channels with dynamic statistical properties. Among time-varying channels, a prominent example is the time-correlated channel [14, 15]. In real-world communication processes, there exist numerous instances of time-correlated channels. For example, user mobility usually leads to time-correlated fading channels in mobile communication [16, 17]. As a consequence of moderate user mobility, channel fading gains become interdependent between consecutive time slots, resulting in the well-known Markov fading channels [18, 19]. Hence, establishing generalization bounds for learning-based communication systems over time-correlated channels is challenging, as statistical learning theory, which significantly depends on the I.I.D. channel assumption, is not applicable in this scenario.
In this paper, we will establish online optimization algorithms [20, 21] for learning communication systems over time-correlated channels with solid theoretical guarantees. The outlined online optimization procedure is depicted in Fig. 1. To be specific, we focus on two types of channels: 1) time-correlated fading channel with additive white Gaussian noise, where the fading distribution is unknown and time-correlated; and 2) time-correlated additive noise channel with an unknown and time-correlated noise distribution. For the time-correlated fading channel, the transceiver is equipped with a fixed codebook (constellation) and is tasked with optimizing the channel decoder using collected channel input-output pairs data. Besides, based on the proposed convex surrogate loss, we can regard such online channel decoder learning problem as an online convex optimization problem. In the case of time-correlated additive noise channels, the goal of transceiver becomes to select optimal codebooks during the whole communication process. We model this online codebook learning problem as a multi-armed bandit problem: the transceiver selects optimal codebooks from a predefined super-codebook using empirical data collected in real-time.
To harness the temporal dependence of considered channels for addressing these two tasks, we separately propose an algorithm based on the optimistic Online Mirror Descent (OMD) framework within online optimization [22, 23]. The key advantage of proposed algorithms lies in their ability to exploit the inherent distribution dependency of benign environments to improve the performance of learned decisions [24]. Then we summarize the main contributions of this work as follows,
  • For the first time, we formulate the task of learning channel decoder or codebook for communication over time-correlated channels as an online optimization problem without relying on the I.I.D. channel assumption, which is more in line with real communication scenarios.
  • We devise various algorithms to tackle the online optimization problem of learning communication systems over time-correlated channels using the optimistic OMD framework. Furthermore, we offer theoretical guarantee for our proposed algorithms.
  • To further support our theoretical framework, we perform simulation experiments to validate the efficacy of proposed methods. Empirical results confirm that our approaches utilize channel correlation to surpass baseline methods, matched with our theoretical discoveries.
The remainder of the paper is structured as follows: In Section II, we present a review of prior related works on learning-based communication systems and online optimization theory. Section III introduces two distinct time-correlated channels and outlines the online optimization problem addressed in this study. Section IV and V detail the development of online optimization algorithms for learning channel decoder and codebook under these two time-correlated channels respectively. Section VI encompasses numerical simulations, while Section VII concludes the paper.
II Related works
In this section, we begin by presenting recent studies that have utilized machine learning techniques in the design of communication systems. Next, we provide an overview of research in online optimization, with an emphasiss on the application of optimistic OMD on predictable benign environments.
As previously mentioned, the remarkable success of machine learning algorithms has spurred interest in their application to optimize communication systems [6]. Some endeavors intend to replace existing components of communication systems with modules learned by empirical data [25]. Another method involves adapting conventional algorithms by integrating deep neural networks [26]. On the other hand, some recent works leverage statistical learning theory for establishing generalization bounds of learned communication systems. For instance, [9] utilize channel input-output pair data to optimize channel decoders and constellations, and derive generalization bounds for learned communication schemes using Rademacher complexity. Besides, [10, 11] employ the probably approximately correct (PAC) learning framework to provide theoretical guarantees for learned communication systems under discrete memoryless channels. Nonetheless, previous works have predominantly conducted a theoretical analysis based on the I.I.D. channel assumption, which is not applicable to with most practical communication scenarios.
Online optimization can effectively model numerous online machine learning problems, encompassing online convex optimization [20, 27], prediction with expert advice [28], multi-armed bandit [29, 30] and more. This framework deals with a sequential decision problem, where a learner repeatedly takes actions within a feasible set and encounters a potentially adversarial loss function from the environment [20, 21]. During this decision-making process, the learner endeavors to devise an algorithm that minimizes regret, defined as the disparity between the total loss incurred by the learner and that of the best decision in hindsight. Departing from statistical learning theory, online optimization theory can furnish theoretical guarantees for algorithms without relying on the I.I.D. assumption. Recently, it has been noted that practical environments often exhibit predictable patterns, which the learner can leverage to achieve reduced regret [22, 24]. To pursue this objective, an optimistic Online Mirror Descent (OMD) framework has been proposed to harness environmental correlations for improving the performance of learned decisions [22, 24]. For instance, [23] employs optimistic OMD to solve the multi-armed bandit problem in predictable environments, while [31] utilizes it to address the stochastically extended adversarial online convex optimization problem.
III Problem Formulation
III-A Notation Conventions
We use standard notation or define it before its first use, and here only focus on main conventions. The Euclidean norm for a vector $\mathbf{v}\in\mathbb{R}^{d}$ is denoted by $\|\mathbf{v}\|_{2}$. The Frobenius norm and spectral norm for a matrix $\mathbf{A}\in\mathbb{R}^{d\times d}$ are denoted by $\|\mathbf{A}\|_{F}:=\sqrt{tr(\mathbf{A}^{T}\mathbf{A})}$ and $\|\mathbf{A}\|_{2}:=\sup\{\|\mathbf{A}\mathbf{v}\|_{2}:\|\mathbf{v}\|_{2}=1\}$ respectively. We define the inner product of two matrices $\mathbf{A}\in\mathbb{R}^{d\times d}$ and $\mathbf{B}\in\mathbb{R}^{d\times d}$ as $\langle\mathbf{A},\mathbf{B}\rangle=tr(\mathbf{A}^{T}\mathbf{B})$. For $n\in\mathbb{N}_{+}$, the set $\{1,2,...,n\}$ is denoted by $[n]$. The cardinality of a finite set $\mathcal{X}$ is denoted by $|\mathcal{X}|$. For a set $\mathcal{X}$, $\mbox{conv}(\mathcal{X})$ means the convex hull of $\mathcal{X}$. The indicator of an event $\mathcal{A}$ is denoted by $\mathbbm{1}\{\mathcal{A}\}$. We denote $\max\{r,0\}$ as $[r]_{+}$, where $r\in\mathbb{R}$. We use $\mathcal{O}(\cdot)$ to hide numerical constants in our upper bound and use $\tilde{\mathcal{O}}(\cdot)$ to additionally hide logarithmic factors.
III-B Time-varying Channel Models
In this paper, we assume that the transmitter sends messages to the receiver over multiple rounds, with the total number of rounds denoted as $T$. We consider the problem of communication over time-varying channels defined below:
$Y_{t}=f_{t}(X_{t})+Z_{t},$ (1)
where $X_{t}\in\mathbb{R}^{d}$ is a codeword which is chosen from a codebook $C_{t}=\{\mathbf{x}_{t}^{j}\}_{j\in[M]}$ with a uniform probability and $Y_{t}\in\mathbb{R}^{d}$ is the corresponding channel output in $t$-th round. $f_{t}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ is a channel transformation, and $Z_{t}\in\mathbb{R}^{d}$ is a channel noise statistically independent of the input $X$ and the transformation $f_{t}$. In this paper, we assume that the codebook $C_{t}\in\mathcal{C}\subset(R^{d})^{M}$ and $\mathcal{C}$ is with the power constraint, i.e., $\mathcal{C}=\{C:\|\mathbf{x}^{j}\|_{2}\leq\gamma_{X},\forall j\in[M]\}$.
The statistical properties of channel transformation $f_{t}$ and channel noise $Z_{t}$ across different rounds are assumed to be different in this work. In other words, we consider time-varying channels and thus the distributions of $f_{t}$ and $Z_{t}$ can vary at each round during the communication process.
In this paper, we delve into two specific cases of the time-varying channel model described earlier, both of which hold high relevance in realistic communication scenarios. The first case involves a time-correlated fading channel, assuming that $f_{t}$ is a time-correlated linear transformation while the channel noise $Z_{t}$ is an independently and identically distributed (I.I.D.) Gaussian noise. The second case pertains a time-correlated additive noise channel, where $f_{t}$ is an identity transformation and $Z_{t}$ is a time-correlated additive channel noise.
III-B1 Time-correlated Fading Channels
In the first considered channel model, we fix the codebook $C_{t}=C=\{\mathbf{x}^{j}\}_{j\in[M]},\forall t\in[T]$ for the transceiver at each round. The channel transformation is assumed to be the linear transformation and it is thus similar to the common fading channel in wireless communications [32, 33]. Hence, Eq. (1) becomes
$Y_{t}=\mathbf{H}_{t}X_{t}+W_{t},$ (2)
where $\mathbf{H}_{t}\in\mathbb{R}^{d\times d}$ is the time-correlated channel gain of this fading channel. The distribution of $\mathbf{H}_{t}$ is denoted by $F_{\mathbf{H}_{t}}$, which is different at each round $t$. Notice that we do not assume any specific distribution for $\mathbf{H}_{t}$, and $\{F_{\mathbf{H}_{t}}\}_{t\in[T]}$ is entirely unknown to the transceiver. $W_{t}\in\mathbb{R}^{d}$ is an I.I.D. additive white Gaussian noise with zero mean and its variance is denoted by $\sigma_{W}^{2}$. The distribution of $W_{t}$ is denoted as $F_{W}$, and $W_{t}$ is independent of the input and the channel gain.
[…]
For this channel, we assume that the decoding rule is chosen from the class of nearest neighbor decoder with a linear kernel operating on the channel output, i.e., given a channel output $\mathbf{y}_{t}^{j}$, the index of the decoded codeword is selected as:
$j_{t}(\mathbf{y}^{j}_{t})\in\mathop{\arg\min}\limits_{j^{\prime}\in[M]}\|\mathbf{x}^{j^{\prime}}-\mathbf{G}_{t}\mathbf{y}^{j}_{t}\|_{2},$ (3)
where $\mathbf{y}_{t}^{j}=\mathbf{H}_{t}\mathbf{x}^{j}+W_{t}$ and $\mathbf{G}_{t}\in\mathbb{R}^{d\times d}$ is a linear kernel operating on the channel output $\{\mathbf{y}^{j}_{t}\}_{j\in[M]}$.
Now we explain this choice of decoding rule. Initially, the selection of the nearest neighbor decoder is based on its optimality under the condition that $\mathbf{H}_{t}=\mathbf{G}_{t}$ and $W$ is Gaussian [32], but we do not assume that the channel satisfies this condition here [34]. Furthermore, the use of the linear kernel $\mathbf{G}_{t}$ on $\mathbf{y}_{t}$ shares similarities with the channel equalizer widely employed in communication systems [3, 32], aiming to approximate the channel gain $\mathbf{H}_{t}$ and minimize its influence on decoding. For simplicity, we refer to the linear kernel $\mathbf{G}_{t}$ as the channel decoder in the subsequent discussions.
We then define the $t$-th round expected error probability of the channel decoder $\mathbf{G}_{t}$ learned in the $t$-th round given that $M$ distinct codewords in $C$ are transmitted:
$\mathbb{P}_{t}(\mathbf{G}_{t}):=\frac{1}{M}\sum_{j=1}^{M}P\Big{\{}\|\mathbf{x}^{j_{t}}-\mathbf{G}_{t}\mathbf{y}_{t}^{j}\|_{2}^{2}<\|\mathbf{x}^{j}-\mathbf{G}_{t}\mathbf{y}_{t}^{j}\|_{2}^{2}\Big{\}},$ (4)
where $j_{t}:=\arg\min_{j^{\prime}\in[M]\setminus j}\|\mathbf{x}^{j^{\prime}}-\mathbf{G}_{t}\mathbf{y}_{t}^{j}\|_{2}^{2}$.
Then, we consider conducting an online optimization procedure to identify the optimal channel decoder $\mathbf{G}_{t}$. The online optimization goal related to this motivation is to leverage channel output samples $\{\mathbf{y}_{t}^{j}\}_{j\in[M]},\forall t\in[T]$ generated at each round to construct the channel decoder $\mathbf{G}_{t}\in\mathcal{G}$ with minimal expected error probability for each round. In this paper, we assume that $\mathcal{G}=\{\mathbf{G}\in\mathbb{R}^{d\times d}:\|\mathbf{G}\|_{F}\leq D\}$.
Given that the transceiver is provided with channel output samples related to the channel gain $\mathbf{H}_{t}$ and the channel noise $W_{t}$ at each round, the empirical symbol error rate of the channel decoder $\mathbf{G}_{t}$ in the $t$-th round can be defined as
$\ell_{t}(\mathbf{G}_{t}):=\frac{1}{M}\sum_{j=1}^{M}\mathbbm{1}\Big{\{}\|\mathbf{x}^{j_{t}}-\mathbf{G}_{t}\mathbf{y}_{t}^{j}\|_{2}^{2}<\|\mathbf{x}^{j}-\mathbf{G}_{t}\mathbf{y}_{t}^{j}\|_{2}^{2}\Big{\}}.$ (5)
Hence, the transceiver can utilize this symbol error rate to carry out the online optimization protocol for learning the channel decoder $\mathbf{G}_{t}$. We design the online convex optimization algorithms to learn $\mathbf{G}_{t}$ in Section IV.
III-B2 Time-correlated Additive Noise Channels
We then focus on the second specific channel model considered in this paper, and introduce the corresponding online optimization problem. For this channel, the channel transformation is assumed to be the identity mapping, i.e., $f_{t}=\mathbf{I}$, so Eq. (1) becomes
$Y_{t}=X_{t}+Z_{t},$ (6)
where $Z_{t}\in\mathbb{R}^{d}$ is a time-correlated channel noise with the distribution $F_{Z_{t}}$, which is statistically independent of the input. Similarly, we do not make any assumptions on the distribution of $Z_{t}$, and $\{F_{Z_{t}}\}_{t\in[T]}$ is completely unknown to the transceiver. We suppose that the channel noise is time-correlated, i.e., $Z_{t}$ is dependent on $\{Z_{\tau}\}_{\tau\in[t-1]}$ from the previous round. We can also consider a first-order Markov noisy channel: $Z_{t+1}=Z_{t}+\boldsymbol{\epsilon}_{t}$, where $\boldsymbol{\epsilon}_{t}\in\mathbb{R}^{d}$ is a random vector. This channel is pertinent to practical communication scenarios, where the presence of thermal noise in the receiver leads to gradual variations.
For this channel, we fix the channel decoder $\mathbf{G}_{t}=I$ for each round since the channel transformation $f_{t}$ is an identity function in this scenario. Similar to the time-correlated fading channel, we also choose the nearest neighbor decoding rule based on the Euclidean distance, i.e., given a channel output $\mathbf{y}_{t}^{j}$, the index of the decoded codeword is chosen as:
$j_{t}(\mathbf{y}^{j}_{t})\in\mathop{\arg\min}\limits_{j^{\prime}\in[M]}\|\mathbf{x}_{t}^{j^{\prime}}-\mathbf{y}^{j}_{t}\|_{2}$ (7)
where $\mathbf{y}_{t}^{j}=\mathbf{x}_{t}^{j}+Z_{t},\forall j\in[M]$ and the codeword $\mathbf{x}_{t}^{j}$ is from the codebook $C_{t}=\{\mathbf{x}_{t}^{j}\}_{j\in[M]}$. Notice that we do not assume that the time-correlated noise is Gaussian and thus this decoding rule is only selected for its simplicity [9, 34].
Analogously, we can define the expected error probability over the $t$-th round channel noise $Z_{t}\sim F_{Z_{t}}$ given that $M$ distinct codewords in $C_{t}$ are transmitted.
$\mathbb{P}_{t}(C_{t}):=\frac{1}{M}\sum_{j=1}^{M}P\Big{\{}\|\mathbf{x}_{t}^{j_{t}}-\mathbf{y}_{t}^{j}\|_{2}^{2}<\|\mathbf{x}_{t}^{j}-\mathbf{y}_{t}^{j}\|_{2}^{2}\Big{\}},$ (8)
where $j_{t}:=\arg\min_{j^{\prime}\in[M]\setminus j}\|\mathbf{x}_{t}^{j^{\prime}}-\mathbf{y}_{t}^{j}\|_{2}^{2}$. $C_{t}=\{\mathbf{x}_{t}^{j}\}_{j\in[M]}$ denotes the codebook chosen in the $t$-th round.
The online optimization goal for this scenario is to utilize the channel output samples $\{\mathbf{y}_{t}^{j}\}_{j\in[M]},\forall t\in[T]$ for selecting the codebook $C_{t}\in\mathcal{C}\subset(R^{d})^{M}$ with minimal expected error probability at each round.
Based on channel output samples related to $Z_{t}$, the symbol error rate of codebook $C_{t}$ can be defined as
$\ell_{t}(C_{t}):=\frac{1}{M}\sum_{j=1}^{M}\mathbbm{1}\Big{\{}\|\mathbf{x}_{t}^{j_{t}}-\mathbf{y}_{t}^{j}\|_{2}^{2}<\|\mathbf{x}^{j}_{t}-\mathbf{y}_{t}^{j}\|_{2}^{2}\Big{\}},$ (9)
Similarly, the transceiver can make use of this symbol error rate as the loss function to perform the online optimization procedure for learning the codebook $C_{t}$. We devise the multi-armed bandit algorithm to learn $C_{t}$ in Section V.
III-C Online Optimization Procedure
As mentioned above, this paper models the problem of learning the channel decoder $\mathbf{G}_{t}$ or codebook $C_{t}=\{\mathbf{x}_{t}^{j}\}_{j\in[M]}$ as an online optimization problem. Assume that there is a transmitter-receiver pair, and the transmitter sends codewords to the receiver over the time-correlated channel. Let $\mathcal{D}$ be a feasible set and $\ell:\mathcal{D}\rightarrow\mathbb{R}$ be a loss function. In general, at round $t\in[T]$, the transceiver carries out the following step:
  1. 1.
    The receiver makes a decision $D_{t}\in\mathcal{D}$, which can be either a channel decoder $\mathbf{G}_{t}$ or a codebook $C_{t}$. Following [9], we also assume that the receiver can transmit the learned codebook $C_{t}$ to the transmitter via the reliable feedback link.
  2. 2.
    The transmitter then sends codewords to the receiver over the time-correlated channel, and then the receiver calculates the symbol error rate based on the received channel output as the loss function $\ell_{t}(D_{t})$.
  3. 3.
    The receiver leverages $\ell_{t}(D_{t})$ to run the online optimization algorithm for making the next decision.
The objective of the considered online optimization problem is to construct a sequence of decisions $\{D_{t}\}_{t\in[T]}$, which minimizes the regret over $T$ rounds, defined as
$\mbox{Reg}_{T}:=\sum_{t=1}^{T}\Big{(}\ell_{t}(D_{t})-\ell_{t}(D^{*})\Big{)}$ (10)
where $D^{*}:=\arg\min_{D\in\mathcal{D}}\sum_{t=1}^{T}\ell_{t}(D)$.
Subsequently, we will show the relationship between minimizing this regret and minimizing the average expected error probability denoted by $\frac{1}{T}\sum_{t=1}^{T}\mathbb{P}_{t}(D_{t})$. Leveraging this insight, we design various algorithms within a general optimistic online mirror descent (OMD) framework [22, 24] to learn channel decoders or codebooks for communication over time-correlated channels, backed by solid theoretical guarantees. The benefit of this framework is to utilize the distribution dependence within predictable environments across different rounds for improving online optimization procedures, which is suitable for time-correlated channels considered in this paper.
Notice that we do not delve into the computational complexity or practical applications of proposed algorithms in this paper. Our focus is to explore theoretical performance limits of communication systems learned by devised algorithms.
IV Learning Channel decoder via Online Convex Optimization
In this section, we consider the time-correlated fading channel and fix the codebook $C_{t}=C=\{\mathbf{x}^{j}\}_{j\in[M]},\forall t\in[T]$ with the constant modulus constraint $\|\mathbf{x}^{j}\|_{2}=\gamma_{X},\forall j\in[M]$ [35] for each round. Hence, we only focus on designing algorithms to learn channel decoders $\{\mathbf{G}_{t}\}_{t\in[T]}$ to minimize the expected error probability defined in Eq. (4). To conserve space, the proofs for all theorems below are provided in the appendix.
IV-A Preliminaries
In this subsection, we first introduce some vital physical quantities and typical assumptions related to the considered channel. Next, we introduce a hinge-type surrogate loss and leverage it to solve the online decoder learning problem.
At first, we define the variance of the time-correlated channel gain $\{\mathbf{H}_{t}\}_{t\in[T]}$ as
$(\sigma_{\mathbf{H}_{t}})^{2}:=\mathbb{E}_{\mathbf{H}_{t}\sim F_{\mathbf{H}_{t}}}\Big{[}\|\mathbf{H}_{t}-\mathbf{U}_{t}\|_{F}^{2}|\mathcal{F}_{t-1}\Big{]},$ (11)
where $\mathbf{U}_{t}:=\mathbb{E}_{\mathbf{H}_{t}\sim F_{\mathbf{H}_{t}}}[\mathbf{H}_{t}|\mathcal{F}_{t-1}]$ denotes the mean matrix of $\mathbf{H}_{t}$ and $\mathcal{F}_{t-1}$ denotes the $\sigma$-algebra generated by $(\mathbf{H}_{1},\mathbf{H}_{2},...,\mathbf{H}_{t-1})$.
Then we define the cumulative variance of $\{\mathbf{H}_{t}\}_{t\in[T]}$ as
$(\sigma^{\mathbf{H}}_{1:T})^{2}:=\mathbb{E}\Big{[}\sum_{t=1}^{T}(\sigma_{\mathbf{H}_{t}})^{2}\Big{]}.$ (12)
Furthermore, we define the cumulative variation of time-correlated channel gain $\{\mathbf{H}_{t}\}_{t\in[T]}$ as
$(\Sigma^{\mathbf{H}}_{1:T})^{2}:=\mathbb{E}\Big{[}\sum_{t=1}^{T}\|\mathbf{U}_{t}-\mathbf{U}_{t-1}\|^{2}_{F}\Big{]},$ (13)
where $\mathbf{U}_{0}=\mathbf{0}$. This quantity reflects the correlation of the channel gain between the previous and current rounds. We observe that $(\Sigma^{\mathbf{H}}_{1:T})^{2}$ decreases when $\{\mathbf{H}_{t}\}_{t\in[T]}$ exhibit high interdependence across different rounds.
Additionally, the following standard assumption widely used in the theoretical analysis of learning-based communication systems [9, 34] is required.
Assumption 1 (Bounded Channel Gain and Channel Noise)
The channel gain is bounded by $\gamma_{H}$ and the additive Gaussian noise is bounded by $\gamma_{W}$ at each round, i.e., for any $t\in[T]$, we have $\mathbb{E}\|\mathbf{H}_{t}\|_{F}\leq\gamma_{H}$ and $\mathbb{E}\|W_{t}\|_{2}=\sigma_{W}\leq\gamma_{W}$.
Based on this assumption, we find that channel outputs $\{\mathbf{y}_{t}^{j}\}_{j\in[M],t\in[T]}$ are bounded by $\sqrt{2(\gamma_{X}\gamma_{H})^{2}+2(\gamma_{W})^{2}}$ and we denote it by $L$ in the following.
We then provide a convex surrogate loss for designing algorithms to learn the channel decoder $\mathbf{G}_{t}$ via upper-bounding the expected error probability in Eq. (4). Notice that we know the inequality $\mathbbm{1}\{t<0\}\leq[r-t]_{+}$ holds, where $r\geq 1$. Hence, the expected error probability can be bounded as follows
$\mathbb{P}_{t}(\mathbf{G}_{t})$ $=\sum_{j=1}^{M}\frac{P\big{\{}\exists j^{\prime}\neq j:\|\mathbf{x}^{j^{\prime}}-\mathbf{G}_{t}\mathbf{y}_{t}^{j}\|^{2}_{2}\leq\|\mathbf{x}^{j}-\mathbf{G}_{t}\mathbf{y}_{t}^{j}\|^{2}_{2}\big{\}}}{M}$ (14)
$\stackrel{{\scriptstyle\text{(a)}}}{{\leq}}\sum_{j=1}^{M}\sum^{M}_{j^{\prime}\neq j}\frac{P\big{\{}\|\mathbf{x}^{j^{\prime}}-\mathbf{G}_{t}\mathbf{y}_{t}^{j}\|^{2}_{2}<\|\mathbf{x}^{j}-\mathbf{G}_{t}\mathbf{y}_{t}^{j}\|^{2}_{2}\big{\}}}{M}$
$\stackrel{{\scriptstyle\text{(b)}}}{{\leq}}\sum_{j=1}^{M}\sum^{M}_{j^{\prime}\neq j}\frac{\mathbb{E}\big{[}r-\|\mathbf{x}^{j^{\prime}}-\mathbf{G}_{t}\mathbf{y}_{t}^{j}\|^{2}_{2}+\|\mathbf{x}^{j}-\mathbf{G}_{t}\mathbf{y}_{t}^{j}\|^{2}_{2}\big{]}_{+}}{M},$
where $r\geq 1$. (a) holds based on the Boole’s inequality, and (b) follows from the fact that $\mathbb{E}[\mathbbm{1}\{\|\mathbf{x}^{j^{\prime}}-\mathbf{G}_{t}\mathbf{y}_{t}^{j}\|^{2}_{2}-\|\mathbf{x}^{j}-\mathbf{G}_{t}\mathbf{y}_{t}^{j}\|^{2}_{2}\leq 0\}]\leq\mathbb{E}\big{[}r-\|\mathbf{x}^{j^{\prime}}-\mathbf{G}_{t}\mathbf{y}_{t}^{j}\|^{2}_{2}+\|\mathbf{x}^{j}-\mathbf{G}_{t}\mathbf{y}_{t}^{j}\|^{2}_{2}\big{]}_{+}$ when $r$ satisfies $r\geq 1$.
Based on the above observation, we utilize the following hinge-type loss as the convex surrogate loss function with respect to $\mathbf{G}_{t}$ with the parameter $r_{t}\geq 1$:
~$\tilde{\ell}_{t}(\mathbf{G}_{t}):=\sum_{j=1}^{M}\sum^{M}_{j^{\prime}\neq j}\frac{\big{[}r_{t}-\|\mathbf{x}^{j^{\prime}}-\mathbf{G}_{t}\mathbf{y}_{t}^{j}\|^{2}_{2}+\|\mathbf{x}^{j}-\mathbf{G}_{t}\mathbf{y}_{t}^{j}\|^{2}_{2}\big{]}_{+}}{M}.$ (15)
The subgradient of this surrogate loss function ~$\tilde{\ell}_{t}(\mathbf{G}_{t})$ is
$\nabla_{\mathbf{G}_{t}}\tilde{\ell}_{t}(\mathbf{G}_{t})$ $=\frac{2}{M}\sum_{j=1}^{M}\sum^{M}_{j^{\prime}\neq j}\mathbbm{1}^{j,j^{\prime}}_{t}(r_{t})(\mathbf{x}^{j^{\prime}}-\mathbf{x}^{j})(\mathbf{y}_{t}^{j})^{T},$ (16)
where $\mathbbm{1}^{j,j^{\prime}}_{t}(r_{t}):=\mathbbm{1}\{\|\mathbf{x}^{j^{\prime}}-\mathbf{G}_{t}\mathbf{y}_{t}^{j}\|^{2}_{2}-\|\mathbf{x}^{j}-\mathbf{G}_{t}\mathbf{y}_{t}^{j}\|^{2}_{2}\leq r_{t}\}$.
Based on the proposed surrogate loss, we can redefine the regret defined in Eq. (10) for this scenario as follows,
$\mbox{Reg}_{T}:=\sum_{t=1}^{T}\Big{(}\tilde{\ell}_{t}(\mathbf{G}_{t})-\tilde{\ell}_{t}(\mathbf{G}^{*})\Big{)},$ (17)
where $\mathbf{G}^{*}:=\arg\min_{\mathbf{G}\in\mathcal{G}}\sum_{t=1}^{T}\tilde{\ell}_{t}(\mathbf{G})$.
IV-B Optimistic OMD with Euclidean Regularizer
Using the proposed hinge-type surrogate loss, we can regard the online channel decoder learning problem as an online convex optimization problem minimizing the regret defined in Eq. (17). In this section, we devise an online optimization algorithm based on the optimistic OMD framework to learn the channel decoder $\mathbf{G}_{t}$ on the fly.
Specifically, during the online optimization process, the transceiver stores two sequences $\{\mathbf{G}_{t}\}_{t=1}^{T}$ and $\{\mathbf{G}^{\prime}_{t}\}_{t=1}^{T}$. At each round $t\in[T]$, the transceiver initially uses a hint matrix $\mathbf{M}_{t}\in\mathcal{G}$, which incorporates specific prior knowledge of the unknown channel gain $\mathbf{H}_{t}$, to construct the channel decoder $\mathbf{G}_{t}$. Then, the transceiver utilizes the learned $\mathbf{G}_{t}$ to finish one round of communication and calculates the corresponding surrogate loss ~$\tilde{\ell}_{t}(\mathbf{G}_{t})$.
Then we introduce the procedure of optimistic OMD [24] below, which is defined as the following two step updates
$\mathbf{G}_{t}$ $=\underset{\mathbf{G}\in\mathcal{G}}{\arg\min}\big{\{}\langle\mathbf{M}_{t},\mathbf{G}\rangle+\mathcal{B}_{\psi_{t}}(\mathbf{G},\mathbf{G}^{\prime}_{t})\big{\}},$ (18)
$\mathbf{G}^{\prime}_{t+1}$ $=\underset{\mathbf{G}\in\mathcal{G}}{\arg\min}\big{\{}\langle\nabla\tilde{\ell}_{t}(\mathbf{G}_{t}),\mathbf{G}\rangle+\mathcal{B}_{\psi_{t}}(\mathbf{G},\mathbf{G}^{\prime}_{t})\big{\}},$
where $\mathcal{B}_{\psi}(X,Y)=\psi(X)-\psi(Y)-\langle\nabla\psi(Y),X-Y\rangle$ denotes the Bregman divergence induced by a differentiable convex function $\psi$, which is called the regularizer. We allow the regularizer $\psi_{t}$ to be time-varying in this paper.
Given that the channel gain $\{\mathbf{H}_{t}\}_{t\in[T]}$ are mutually dependent across different rounds in this scenario, we set the hint matrix $\mathbf{M}_{t}$ as ~$\nabla\tilde{\ell}_{t-1}(\mathbf{G}_{t-1})$, i.e., the last-round gradient, to enhance the online optimization process. In addition, we can set $\mathbf{G}_{1}=\mathbf{G}^{\prime}_{1}$ to be an arbitrary matrix in $\mathcal{G}$.
In this section, we set the regularizer as the Euclidean norm [31], i.e., $\psi_{t}(\mathbf{G})=\frac{1}{2\eta_{t}}\|\mathbf{G}\|^{2}_{F}$ with the learning rate
$\eta_{t}=\frac{D}{\sqrt{1+\sum_{\tau=1}^{t-1}\|\nabla\tilde{\ell}_{\tau}(\mathbf{G}_{\tau})-\mathbf{M}_{\tau}\|_{F}^{2}}}.$ (19)
To sum up, the update rules in Eq. (18) become
$\mathbf{G}_{t}$ $=\Pi_{\mathcal{G}}[\mathbf{G}^{\prime}_{t}-\eta_{t}\nabla\tilde{\ell}_{t-1}(\mathbf{G}_{t-1})],$ (20)
$\mathbf{G}^{\prime}_{t+1}$ $=\Pi_{\mathcal{G}}[\mathbf{G}^{\prime}_{t}-\eta_{t}\nabla\tilde{\ell}_{t}(\mathbf{G}_{t})],$
where $\Pi_{\mathcal{G}}$ denotes the Euclidean projection onto the feasible domain $\mathcal{G}$. In fact, the proposed approach performs gradient descent twice at each round. Besides, the step size $\{\eta_{t}\}_{t\in[T]}$ is chosen adaptively, similar to self-confident tuning [36].
Input: The number of communication round $T$, step size $\{\eta_{t}\}_{t\in[T]}$, the parameters $\{r_{t}\}_{t\in[T]}$ in surrogate loss.
1 Initialize: $\mathbf{G}_{1}\in\mathcal{G}$, and $\mathbf{G}_{1}=\mathbf{G}^{\prime}_{1}$ ;
2 for round $t\in[T]$  do
3       Update the channel decoder: $\mathbf{G}_{t}=\Pi_{\mathcal{G}}[\mathbf{G}^{\prime}_{t}-\eta_{t}\nabla\tilde{\ell}_{t-1}(\mathbf{G}_{t-1})]$;
4       The transmitter sends codewords to the receiver;
5       The receiver calculates ~$\tilde{\ell}_{t}(\mathbf{G}_{t})$;
6       Update the auxiliary channel decoder: $\mathbf{G}^{\prime}_{t+1}=\Pi_{\mathcal{G}}[\mathbf{G}^{\prime}_{t}-\eta_{t}\nabla\tilde{\ell}_{t}(\mathbf{G}_{t})]$;
7
8 end for
Algorithm 1 Optimistic OMD for Learning Channel decoder
The protocol of the proposed algorithm is illustrated in Algorithm 1. In the following, we establish the theoretical guarantee of the proposed algorithm via offering an upper bound on the expected error probability with learned channel decoders $\{\mathbf{G}_{t}\}_{t\in[T]}$.
Theorem 1
Under Assumptions 1, if we select the parameters $r_{t}=2d^{*}DL+\frac{1}{\sqrt{T}}$ with the maximum codewords distance $d^{*}:=\max_{i\neq j}\|\mathbf{x}^{i}-\mathbf{x}^{j}\|_{2}\geq\frac{1}{2DL}$, we have
$\frac{1}{T}\sum_{t\in[T]}\mathbb{P}(\mathbf{G}_{t})-\frac{1}{T}\sum_{t\in[T]}\mathbb{E}[\tilde{\ell}_{t}(\mathbf{G}^{*})]$ (21)
$=\mathcal{O}\Big{(}\frac{1}{T}\sqrt{M\sum_{t\in[T]}\sum_{j\in[M]}\mathbb{E}\|\mathbf{y}_{t}^{j}-\mathbf{y}_{t-1}^{j}\|_{2}^{2}}\Big{)}.$
Remark 1
Theorem 1 provides a performance guarantee of the learned channel decoder $\{\mathbf{G}_{t}\}_{t\in[T]}$: if the considered online learning problem is realizable, i.e., $\exists\mathbf{G}^{*}\in\mathcal{G}$ such that the term $\sum_{t\in[T]}\mathbb{E}[\tilde{\ell}_{t}(\mathbf{G}^{*})]$ equals to zero [20, 28], the average expected error probability satisfies $\frac{1}{T}\sum_{t\in[T]}\mathbb{P}(\mathbf{G}_{t})\rightarrow 0$ as $T\rightarrow\infty$. Besides, the term $\|\mathbf{y}_{t}^{j}-\mathbf{y}_{t-1}^{j}\|_{2}^{2}$ reflects the variation of channel outputs across successive rounds transmitted by the same codeword. We observe that as this term decreases, the derived upper bound becomes tighter, indicating that our approach leverages the distribution dependence of $\{\mathbf{H}_{t}\}_{t\in[T]}$ to enhance the efficacy of $\mathbf{G}_{t}$.
To better understand the result of Theorem 1, we utilize the physical quantities introduced before to provide the below corollary based on Theorem 1.
Corollary 1
Under the conditions of Theorem 1, we have
$\frac{1}{T}\sum_{t\in[T]}\mathbb{P}(\mathbf{G}_{t})-\frac{1}{T}\sum_{t\in[T]}\mathbb{E}[\tilde{\ell}_{t}(\mathbf{G}^{*})]$ (22)
$=\mathcal{O}\Big{(}\frac{M}{T}\sqrt{(\sigma^{\mathbf{H}}_{1:T})^{2}}+\frac{M}{T}\sqrt{(\Sigma^{\mathbf{H}}_{1:T})^{2}}+\frac{M\sqrt{T}}{T}\sqrt{\sigma_{W}^{2}}\Big{)},$
where $(\sigma^{\mathbf{H}}_{1:T})^{2}$ follows Eq. (12) and $(\Sigma^{\mathbf{H}}_{1:T})^{2}$ follows Eq. (13).
Remark 2
Corollary 1 implies that the variance $\sigma^{2}_{W}$ of the channel noise $W$, the cumulative variance $(\sigma^{\mathbf{H}}_{1:T})^{2}$ and the cumulative variation $(\Sigma^{\mathbf{H}}_{1:T})^{2}$ of channel gain $\{\mathbf{H}_{t}\}_{t\in[T]}$ can deteriorate the performance of the learned channel decoder $\{\mathbf{G}_{t}\}_{t\in[T]}$. We observe that smaller $(\Sigma^{\mathbf{H}}_{1:T})^{2}$ leads to a tighter upper bound of average expected error probability $\frac{1}{T}\sum_{t\in[T]}\mathbb{P}(\mathbf{G}_{t})$, reflecting that the proposed method utilizes the distribution dependence of the time-correlated channel gain $\{\mathbf{H}_{t}\}_{t\in[T]}$ to enhance the online optimization process.
IV-C Extension: Learning to Communicate over Independently and Identically Distributed Fading Channel
In this subsection, we focus on the I.I.D. fading channel, that is, the channel gains $\mathbf{H}_{1},\mathbf{H}_{2},...,\mathbf{H}_{t}$ are independently sampled from the distribution $F_{\mathbf{H}}$. Consequently, the variation of $\{\mathbf{H}_{t}\}_{t\in[T]}$ between different rounds satisfies $(\Sigma^{\mathbf{H}}_{1:T})^{2}=0$ and the variance of $\mathbf{H}_{t}$ satisfies $(\sigma_{\mathbf{H}_{t}})^{2}=\sigma_{\mathbf{H}}^{2},\forall t\in[T]$.
Hence, the online optimization problem considered before is now converted into a stochastic optimization problem:
$\min_{\mathbf{G}\in\mathcal{G}}$~$\Big{\{}\tilde{\ell}(\mathbf{G})$ (23)
$:=\frac{1}{M}\sum_{j=1}^{M}\sum^{M}_{j^{\prime}\neq j}\big{[}r-\|\mathbf{x}^{j^{\prime}}-\mathbf{G}\mathbf{y}^{j}\|^{2}_{2}+\|\mathbf{x}^{j}-\mathbf{G}\mathbf{y}^{j}\|^{2}_{2}\big{]}_{+}\Big{\}},$
where $r>1$ and $\mathbf{y}^{j}:=\mathbf{H}\mathbf{x}^{j}+W,\forall j\in[M]$. $\mathbf{H}\sim F_{\mathbf{H}}$ is an I.I.D. channel gain with unknown distribution, and $W$ is an I.I.D. additive white Gaussian noise.
Next, we employ the well-established online-to-batch conversion method [20, 21] to solve the problem defined in Eq. (23). Furthermore, we can undertake the generalization analysis of the channel decoder $\bar{\mathbf{G}}$ learned by this way for I.I.D. fading channels. The fundamental idea of this approach is to utilize the I.I.D. channel data in a sequential manner, and execute the online optimization algorithm to minimize the regret. Upon completion of this sequential learning process, the average decision serves as the ultimate output of this online-to-batch conversion procedure.
Input: An online optimization algorithm $\mathcal{A}$
1 for round $t\in[T]$  do
2       Let $\mathbf{G}_{t}$ be the output of algorithm $\mathcal{A}$ for this round;
3       Feed algorithm $\mathcal{A}$ with the loss function ~$\tilde{\ell}_{t}(\mathbf{G}_{t})$ calculated by data $(\mathbf{H}_{t},W_{t})$;
4
5 end for
Return: $\bar{\mathbf{G}}=\frac{1}{T}\sum_{t\in[T]}\mathbf{G}_{t}$
Algorithm 2 Online-to-Batch Conversion
The online-to-batch conversion is performed as illustrated in Algorithm 2. If the online optimization algorithm $\mathcal{A}$ is selected as the Algorithm 1, we can derive the expected error probability of $\bar{\mathbf{G}}$ constructed by $\mathcal{A}$ below.
Theorem 2
If the loss function ~$\tilde{\ell}(\bar{\mathbf{G}})$ is convex w.r.t. $\bar{\mathbf{G}}$, the channel gain $\{\mathbf{H}_{t}\}_{t\in[T]}$ are i.i.d. sampled from $F_{\mathbf{H}}$ and the channel noise $\{W_{t}\}_{t\in[T]}$ are i.i.d. sampled from $F_{W}$, the expected error probability of $\bar{\mathbf{G}}$ satisfies
$\mathbb{P}(\bar{\mathbf{G}})-\min_{\mathbf{G}\in\mathcal{G}}\mathbb{E}_{\mathbf{H},W}[\tilde{\ell}(\mathbf{G})]=\mathcal{O}\Big{(}M\sqrt{\frac{\sigma_{\mathbf{H}}^{2}}{T}}+M\sqrt{\frac{\sigma_{W}^{2}}{T}}\Big{)},$ (24)
where $\mathbb{E}_{\mathbf{H},W}[\tilde{\ell}(\mathbf{G})]$ is the expected risk.
Remark 3
Theorem 2 implies that the expected error probability of $\bar{\mathbf{G}}$ only depends on the variance $\sigma_{\mathbf{H}}^{2}$ of channel gain $\mathbf{H}$ and the variance $\sigma_{W}^{2}$ of channel noise $W$, since $(\Sigma^{\mathbf{H}}_{1:T})^{2}$ reflecting the variation of $\{\mathbf{H}_{t}\}_{t\in[T]}$ becomes zero when channel gain is I.I.D. with the distribution $F_{\mathbf{H}}$. Similarly, if the statistical learning problem defined in Eq. (23) is realizable, i.e., $\exists\mathbf{G}^{*}\in\mathcal{G}$ such that $\mathbb{E}_{\mathbf{H},W}[\tilde{\ell}(\mathbf{G}^{*})]$ equals to zero [12, 13], we have $\mathbb{P}(\bar{\mathbf{G}})\rightarrow 0$, as $T\rightarrow\infty$, providing the theoretical performance guarantee for the proposed online-to-batch conversion method. Hence, $T$ plays a role akin to sample complexity within statistical learning theory [13].
V Learning codebook via Multi-armed Bandit
In this section, we focus on the time-correlated additive noise channel and set the channel decoder $\{\mathbf{G}_{t}\}_{t\in[T]}$ as the identity matrix $\mathbf{I}$. Hence, we only consider devising algorithms to learn the codebook $\{C_{t}\}_{t\in[T]}$ to minimize the expected error probability defined in Eq. (8). Similarly, for brevity, the proofs for all theorems below are available in the appendix.
V-A Preliminaries
Generally, identifying an optimal codebook becomes challenging when the noise distribution is unknown [9, 11]. Even more complex is the dynamic adjustment of the codebook to accommodate channel noise with the time-varying distribution.
One potential approach to tackle this challenge involves selecting codebooks for practical transmission from a pre-defined super-codebook $\mathbf{C}=\{C_{i}\}_{i=1}^{N}$ comprising $|\mathbf{C}|=N$ codebooks [9, 37]. This super-codebook can be statically constructed (such as a grid or a lattice [38]) in advance. To adapt to the changing statistical property of the channel noise, we dynamically choose one of these codebooks used for transmitting codewords in each round. Specifically, we select the codebook based on the symbol error rate of codebooks over the channel noise $Z_{t},\forall t\in[T]$. This approach draws inspiration from the Gibbs-algorithm-based codebook expurgation proposed in [9], and is amenable to practical deployment following the paradigm of Adaptive Modulation and Coding (AMC) widely deployed in communication systems [32, 33].
In this paper, we model this online codebook learning problem as a multi-armed bandit problem [28]. We consider an iterative process spanning $T$ rounds below. In each round, the transceiver selects a codebook $C_{t}$ from the super-codebook $\mathbf{C}$ and uses it for transmitting codewords over $Z_{t}$. Then the receiver calculates the corresponding symbol error rate of this chosen codebook $C_{t}$ as the loss. We denote codewords in $C_{t}$ as $\{\mathbf{x}_{t}^{j}\}_{j\in[M]}$, and define the loss of $C_{t}$ as
$\ell_{t}(C_{t}):=\frac{1}{M}\sum_{j\in[M]}\mathbbm{1}\Big{\{}\|\mathbf{x}_{t}^{j_{t}}-\mathbf{y}_{t}^{j}\|_{2}^{2}\leq\|\mathbf{x}_{t}^{j}-\mathbf{y}_{t}^{j}\|_{2}^{2}\Big{\}},$ (25)
where $\mathbf{y}_{t}^{j}=\mathbf{x}^{j}_{t}+Z_{t}$ and $j_{t}=\arg\min_{j^{\prime}\in[M]\setminus j}\|\mathbf{x}_{t}^{j^{\prime}}-\mathbf{y}_{t}^{j}\|_{2}^{2}$.
Formally, the transceiver selects a binary vector $\mathbf{a}_{t}$ called the index vector from the feasible set $\mathcal{X}:=\{\mathbf{e}_{1},\mathbf{e}_{2},...,\mathbf{e}_{N}\}$, where $\mathbf{e}_{i}$ denotes the $i$-th standard basis vector. In other words, in each round, the transceiver chooses the index $i_{t}\in[N]$ (corresponding to $\mathbf{a}_{t}=\mathbf{e}_{i_{t}}$) of the codebook $C_{t}$. The transceiver then uses $C_{t}$ for transmitting codewords, and suffers loss denoted by $\mathbf{a}_{t}^{T}\boldsymbol{\ell}_{t}=\ell_{t}(C_{t})$, where $\boldsymbol{\ell}_{t}\in[0,1]^{N}$ is a vector including all the symbol error rates of codebooks in $\mathbf{C}$ transmitted over $Z_{t}$. The regret now can be redefined as
$\mbox{Reg}_{T}:=\sum_{t\in[T]}\left[\mathbf{a}_{t}^{T}\boldsymbol{\ell}_{t}-(a^{*})^{T}\boldsymbol{\ell}_{t}\right]=\sum_{t\in[T]}\left[\ell_{t}(C_{t})-\ell_{t}(C^{*})\right],$ (26)
where $\mathbf{a}^{*}:=\min_{\mathbf{a}\in\mathcal{X}}\sum_{t\in[T]}\mathbf{a}^{T}\boldsymbol{\ell}_{t}$ and $C^{*}\in\mathbf{C}$ denotes the codebook corresponding to the index vector $\mathbf{a}^{*}$.
V-B Optimistic OMD with Log-Barrier Regularizer
In this section, we also use the optimistic OMD framework to design an online optimization algorithm for solving the problem of selecting codebooks to communicate over time-correlated additive noise channels. Similarly, the proposed algorithm offers the advantage of utilizing the distribution dependence of such channels for boosting bandit learning processes, suited for the channel considered in this section.
The OMD framework employed in bandit operates on the set $\Omega=\mbox{conv}(\mathcal{X}):=\{\sum_{i\in[N]}\beta_{i}\mathbf{e}_{i}:\sum_{i\in[N]}\beta_{i}=1,\beta_{i}\geq 0,\forall i\in[N]\}$. The update rule of OMD for bandit is $\mathbf{w}_{t}=\arg\min_{\mathbf{w}\in\Omega}\{\langle\mathbf{w},\hat{\boldsymbol{\ell}}_{t-1}\rangle+\mathcal{B}_{\psi}(\mathbf{w},\mathbf{w}_{t-1})\}$ for the regularizer $\psi$ and an unbiased estimator ^$\hat{\boldsymbol{\ell}}_{t-1}$ of the true loss $\boldsymbol{\ell}_{t-1}$. The transceiver then selects the index vector $\mathbf{a}_{t}$ randomly such that $\mathbb{E}[\mathbf{a}_{t}]=\mathbf{w}_{t}$, which corresponds to the codebook $C_{t}$. In essence, $\mathbf{a}_{t}$ is sampled from the probability distribution $\mathbf{w}_{t}$. Then we construct the next ^$\hat{\boldsymbol{\ell}}_{t}$ based on the feedback.
In this section, the optimistic OMD framework also involves maintaining a sequence of auxiliary action $\mathbf{w}^{\prime}_{t}$ updated by ^$\hat{\boldsymbol{\ell}}_{t}$. As mentioned above, Optimistic OMD makes a decision $\mathbf{a}_{t}\sim\mathbf{w}_{t}$ randomly, and $\mathbf{w}_{t}$ is now updated by minimizing $\mathbf{m}_{t}\in[0,1]^{N}$, an optimistic hint of the true loss $\boldsymbol{\ell}_{t}$. Hence, the update rules of optimistic OMD for this scenario become
$\mathbf{w}_{t}$ $=\mathop{\arg\min}\limits_{\mathbf{w}\in\Omega}\big{\{}\langle\mathbf{w},\mathbf{m}_{t}\rangle+\mathcal{B}_{\psi_{t}}(\mathbf{w},\mathbf{w}^{\prime}_{t})\big{\}},$ (27)
$\mathbf{w}^{\prime}_{t+1}$ $=\mathop{\arg\min}\limits_{\mathbf{w}\in\Omega}\big{\{}\langle\mathbf{w},\hat{\boldsymbol{\ell}}_{t}\rangle+\mathcal{B}_{\psi_{t}}(\mathbf{w},\mathbf{w}^{\prime}_{t})\big{\}}.$
Following [23], we set the regularizer as the log-barrier $\psi_{t}(\mathbf{w})=\sum_{i\in[N]}\frac{1}{\eta_{t}}\ln\frac{1}{\mathbf{w}_{i}}$ with learning rate $\eta_{t}$ for deriving our theoretical results. Recall that we consider the time-correlated additive noise channel in this section, i.e., $Z_{t}$ depends on $\{Z_{\tau}\}_{\tau\in[t-1]}$ from previous rounds. Therefore, for utilizing such dependence to enhance the bandit learning process, we set the $i$-th component $\mathbf{m}_{t,i}$ of $\mathbf{m}_{t}$ to be the most recent observed loss of codebook $i\in[N]$. Specifically, $\mathbf{m}_{t,i}$ is set as $\mathbf{m}_{t,i}=\boldsymbol{\ell}_{\alpha_{i}(t),i}$, where $\alpha_{i}(t)$ is defined to be the most recent time when codebook $i$ is chosen prior to round $t$, that is $\alpha_{i}(t):=\max\{\tau<t:i_{\tau}=i\}$ (or 00 if the set is empty).
Input: The number of communication round $T$, step size $\{\eta_{t}\}_{t\in[T]}$.
1 Initialize: $\mathbf{w}^{\prime}_{1}=\arg\min_{\mathbf{w}\in\Omega}\psi_{1}(\mathbf{w})$ ;
2 for round $t\in[T]$  do
3       Update the action: $\mathbf{w}_{t}=\mathop{\arg\min}\limits_{w\in\Omega}\{\langle\mathbf{w},\mathbf{m}_{t}\rangle+\mathcal{B}_{\psi_{t}}(\mathbf{w},\mathbf{w}^{\prime}_{t})\}$;
4       The transmitter sends codewords to the receiver based on the codebook $C_{t}$ corresponding to the index vector $\mathbf{a}_{t}\sim\mathbf{w}_{t}$;
5       The receiver calculates $\mathbf{a}_{t}^{T}\boldsymbol{\ell}_{t}=\ell_{t}(C_{t})$ and constructs the unbiased estimator ^$\hat{\boldsymbol{\ell}}_{t}$ of $\boldsymbol{\ell}_{t}$;
6       Update the auxiliary action: $\mathbf{w}^{\prime}_{t+1}=\mathop{\arg\min}\limits_{\mathbf{w}\in\Omega}\{\langle\mathbf{w},\hat{\boldsymbol{\ell}}_{t}\rangle+\mathcal{B}_{\psi_{t}}(\mathbf{w},\mathbf{w}_{t}^{\prime})\}$;
7
8 end for
Algorithm 3 Optimistic OMD for Learning codebook
The protocol of the presented algorithm for solving the online codebook learning problem under the time-correlated additive noise channel is illustrated in Algorithm 3.
Analogously, we provide the theoretical guarantee for the proposed method via deriving an upper bound on the averaged expected error probability of learned codebooks $\{C_{t}\}_{t\in[T]}$.
Theorem 3
Let ^$\hat{\boldsymbol{\ell}}_{t}$ be an estimator of $\boldsymbol{\ell}_{t}$, satisfying
$\forall i\in[N],\quad\hat{\boldsymbol{\ell}}_{t,i}=\frac{\boldsymbol{\ell}_{t,i}-\mathbf{m}_{t,i}}{\mathbf{w}_{t,i}}\cdot\mathbbm{1}\{i_{t}=i\}+\mathbf{m}_{t,i},$
and set the learning rate $\eta_{t}\leq\frac{1}{162}$ using the doubling trick [23, 28], we have
$\frac{1}{T}\sum_{t\in[T]}\mathbb{P}(C_{t})-\frac{1}{T}\sum_{t\in[T]}\mathbb{E}[\ell_{t}(C^{*})]$ (28)
$=\tilde{\mathcal{O}}\Big{(}\frac{1}{T}\sqrt{\sum_{i\in[N]}\sum_{t\in[T]}\frac{1}{M}\sum_{j\in[M]}\big{|}\mathbb{P}^{j}_{t}(i)-\mathbb{P}^{j}_{t-1}(i)\big{|}}$
$\quad+\frac{1}{T}\sqrt{\sum_{i\in[N]}\sum_{t\in[T]}\frac{1}{M}\sum_{j\in[M]}\sigma[\mathbbm{1}_{t}^{j}(i)]}\Big{)},$
where $\mathbbm{1}_{t}^{j}(i):=\mathbbm{1}\{\|\mathbf{x}_{t}^{j_{t}}-\mathbf{y}_{t}^{j}\|_{2}\leq\|\mathbf{x}_{t}^{j}-\mathbf{y}_{t}^{j}\|_{2}|\mathbf{a}_{t}=\mathbf{e}_{i}\}$ denotes the indicator of misclassifying the $j$-th codeword when the $i$-th codebook is selected in round $t$. $\mathbb{P}^{j}_{t}(i):=\mathbb{E}[\mathbbm{1}_{t}^{j}(i)]$ denotes the expectation of $\mathbbm{1}_{t}^{j}(i)$, and $\sigma[\mathbbm{1}_{t}^{j}(i)]:=\sqrt{\mathbb{E}[\mathbbm{1}_{t}^{j}(i)-\mathbb{P}^{j}_{t}(i)]^{2}}$ denotes the standard deviation of $\mathbbm{1}_{t}^{j}(i)$.
Remark 4
Theorem 3 presents the theoretical performance guarantee of the learned codebook $\{C_{t}\}_{t\in[T]}$: if $\exists C^{*}\in\mathbf{C}$ such that $\sum_{t\in[T]}\ell_{t}(C^{*})$ is sub-linear, i.e., $\sum_{t\in[T]}\ell_{t}(C^{*})=\mathcal{O}(T^{\alpha}),\alpha<1$, then we have $\frac{1}{T}\sum_{t\in[T]}\mathbb{P}(C_{t})\rightarrow 0$, as $T\rightarrow\infty$. In addition, this theorem demonstrates that the performance of $C_{t}$ becomes better if the expected error probability $\mathbb{P}_{t}^{j}(i)$ of the $j$-th codebword in the same codebook $C_{i},\forall i\in[N]$ varies slowly between successive rounds, i.e., the difference $\big{|}\mathbb{P}^{j}_{t}(i)-\mathbb{P}^{j}_{t-1}(i)\big{|}$ is small. This finding reflects that we can learn the optimal codebook $C_{t}$ if the statistical property of $Z_{t}$ does not change heavily, which indicates that the proposed method leverages the distribution dependence of $\{Z_{t}\}_{t\in[T]}$ to improve the bandit learning process.
V-C Case study: $2$-ary codebook for Gaussian channels
To better understand the result from Theorem 3, we consider the below example about utilizing $2$-ary codebook to transmit codewords over Gaussian channels. Specifically, we assume that the time-correlated channel noise $Z_{t}$ is a zero-mean Gaussian noise with the variance of $\sigma^{2}_{Z_{t}}$. Additionally, we consider the $2$-ary codebook below, i.e., we set the number $M$ of codewords satisfies $M=2$ and thus $C_{i}=\{\mathbf{x}^{1}_{i},\mathbf{x}^{2}_{i}\},\forall i\in[N]$. Correspondingly, we denote the distance between the two codewords in the $i$-th codebook as $d_{i}:=\|\mathbf{x}^{1}_{i}-\mathbf{x}^{2}_{i}\|_{2}$.
Then we can directly calculate the expected error probability of the $i$-th codebook over $Z_{t}$ as
[…]
Based on it, we can derive the average expected error probability of the $2$-ary codebook learned by our method for this time-correlated Gaussian channel as follows.
Corollary 2
Under the conditions of Theorem 3, if for any $t\in[T]$, $Z_{t}$ is a zero-mean Gaussian noise with the variance of $\sigma^{2}_{Z_{t}}$, and the number of codewords in any codebooks in $\mathbf{C}$ is set as two, i.e., $M=2$, we have
$\frac{1}{T}\sum_{t\in[T]}\mathbb{P}(C_{t})-\frac{1}{T}\sum_{t\in[T]}\mathbb{E}[\ell_{t}(C^{*})]$ (30)
$=\tilde{\mathcal{O}}\Big{(}\frac{1}{T}\sqrt{\sum_{i\in[N]}\sum_{t\in[T]}\sqrt{Q\big{(}\frac{d_{i}}{2\sigma_{Z_{t}}}\big{)}\Big{(}1-Q\big{(}\frac{d_{i}}{2\sigma_{Z_{t}}}\big{)}\Big{)}}}$
$\quad+\frac{1}{T}\sqrt{\sum_{i\in[N]}\sum_{t\in[T]}\big{|}\sigma_{Z_{t}}-\sigma_{Z_{t-1}}\big{|}}\Big{)}.$
Remark 5
Corollary 2 indicates that smaller difference $|\sigma_{Z_{t}}-\sigma_{Z_{t-1}}|$ of the standard deviation $\{\sigma_{Z_{t}}\}_{t\in[T]}$ between successive rounds makes the performance of the learned codebook $\{C_{t}\}_{t\in[T]}$ better, implying that the proposed method fully leverages the distribution dependence of time-correlated Gaussian noise $\{Z_{t}\}_{t\in[T]}$ to boost the bandit learning process.
(a) Baseline methods for GMD
(b) Baseline methods for LMD
(c) Channel variation of GMD
(d) Channel variation of LMD
Figure 2: For the online decoder learning task, we compare the proposed Euclidean-regularized optimistic OMD with different baseline methods, and show the effect of channel variation on the performance of optimistic OMD.
(a) Baseline methods for GMD
(b) Baseline methods for LMD
(c) Channel variation of GMD
(d) Channel variation of LMD
Figure 3: For the online codebook learning task, we compare the proposed Log-barrier-regularized optimistic OMD with different baseline methods, and show the effect of channel variation on the performance of optimistic OMD.
VI Simulation results
In this section, we conduct simulation experiments to verify the empirical performance of proposed algorithms for the tasks of online decoder learning and online codebook learning, respectively. For the former, we assume that the transceiver is provided with a fixed codebook $C$ of $M$ codewords in $\mathbb{R}^{d}$, which intends to learn channel decoders $\{\mathbf{G}_{t}\}_{t\in[T]}$ for communication $T$ rounds over the time-correlated fading channel. As for the latter, the transceiver aims to select optimal codebooks $\{C_{t}\}_{t\in[T]}$ from a pre-defined super-codebook $\mathbf{C}$ with $N$ codebooks for communication $T$ rounds over the time-correlated additive noise channel. We utilize the average symbol error rate $\frac{1}{t}\sum_{\tau=1}^{t}\ell_{\tau}(D_{\tau}),t\in[T]$ defined in Eq. (5) or Eq. (9) to evaluate the performance of various methods, where $D_{\tau}$ denotes $\mathbf{G}_{\tau}$ or $C_{\tau}$. We run all experiments three times independently with different random seeds, and report the mean and standard deviation of results.
[…]
In the following, we compare the proposed algorithms with various baselines, and explore the effect of channel variation on the proposed methods for the two tasks. We set the code length $d$ and the number $M$ of codewords as $4$ below. The number $K$ of components in mixture distribution is set to $3$.
Initially, we focus on the online decoder learning task. For each round, we use a fixed codebook $C$ with the constant modulus constraint for transmission, and set the Signal-to-Noise Ratio to $24$dB. In addition, we set the degree $\rho$ of channel variation to $0.1$ for two channel distributions. For this task, we compare the proposed method with three baseline methods. The first one is to set the decoder $\mathbf{G}_{t}$ as the identity matrix $\mathbf{I}$, indicating not equalizing the channel output. The second one is to utilize the least squares estimation [32, 33] to construct $\mathbf{G}_{t}$ based on $M$ codewords $\{\mathbf{y}_{t}^{j}\}_{j\in[M]}$ received in each round. We choose Online Gradient Descent (OGD) as the third baseline method, which is widely applied in online convex optimization [20, 21]. To investigate the effect of the channel variation on the proposed method, we vary the degree $\rho$ across $\{1,3,5,10\}$ for both distributions. As illustrated in Fig. 2 (a) and (b), for two channel distributions, the proposed method shows a lower average symbol error rate compared to other baselines, and exhibits reduced variance against random channels. These results imply that the proposed method leverages distribution dependency to achieve superior performance. Moreover, from the results presented in Fig. 2 (c) and (d), the performance of optimistic OMD improves as $\rho$ decreases, suggesting that the distribution dependency assists the proposed method to learn decoders over time-correlated fading channels, matched with our theoretical findings.
(a) $M$ for GMD
(b) $M$ for LMD
Figure 4: For the online decoder learning task, we show the effect of the number $M$ of codewords on the performance of optimistic OMD.
Next, we concentrate on the online codebook learning task over the time-correlated additive noise channel. The pre-constructed super-codebook $\mathbf{C}$ comprises randomly generated codebooks whose codewords are drawn from a uniform distribution in an element-wise manner. The number $N=|\mathbf{C}|$ of these codebooks is fixed as $100$. For this task, we maintain the degree $\rho$ of channel variation at $0.1$ for two distributions. Our method is compared with three baseline methods. The first baseline is to select a fixed codebook from $\mathbf{C}$ for transmission throughout the entire communication process, while the second method entails randomly selecting a codebook from $\mathbf{C}$ in each round. Additionally, we employ a classical multi-armed bandit algorithm known as Exponential-weight for Exploration and Exploitation (EXP3) [20, 28] as another baseline. To examine the impact of channel variation on our proposed method, we set the degree $\rho$ of channel variation across $\{0.5,1,3,5\}$ for conducting simulations. All experimental results are showcased in Fig. 3. In Fig. 3 (a) and (b), the proposed method demonstrates the lowest average symbol error rate over the three other baselines across the two distributions. Notably, EXP3 exhibits a similar performance with the manner of selecting a fixed codebook under Gaussian mixture distribution, and shows considerable variance under Laplace mixture distribution. It means that EXP3 overlooks utilizing the distribution dependency to boost bandit learning processes. Besides, as $\rho$ increases, signifying a more pronounced effect of channel variation, our method yields a higher average error rate, indicating that this method in fact leverages the mild environment dynamics to select the optimal codebook. These empirical observations consistently support and validate our theoretical discoveries.
We proceed to investigate the effect of the number $M$ of codewords on the performance of the proposed method for the online decoder learning task. The experimental setup is the same as the previous simulations. We vary $M$ over $\{4,8,16,32\}$, and the results for two channel distributions are shown in Fig. 4. In Fig. 4, we find that as $M$ decreases, the average error rate of the learned decoder diminishes. This observation aligns with our theoretical results for online decoder learning derived in Section IV: a smaller $M$ tightens the upper bound of expected error probability, thereby improving the performance of learned decoders.
VII Conclusion
In this paper, we consider the problem of learning for communication over time-correlated channels via online optimization. To tackle this challenge, we employ the optimistic OMD framework to develop algorithms for designing communication systems. Furthermore, we provide theoretical guarantees for our methods by deriving sub-linear regret bounds on the expected error probability of learned communication schemes. Our theoretical findings confirm that devised algorithms utilize the distribution dependency of time-correlated channels to improve the performance of learned decoders and codebooks. To verify the effectiveness of our approaches, we conduct extensive simulation experiments and confirm the proposed methods’ superiority over other baselines, thus aligning with our theoretical discoveries.
-A Proof of Theorem 1
Lemma 1 (Proposition 18 in [22])
Let $\Omega$ be a convex compact set, $\psi$ be a convex function on $\Omega$ and $f^{\prime}_{t-1}\in\Omega$. If $f^{\ast}=\mathop{\arg\min}_{f\in\Omega}\left\{\langle x,f\rangle+\mathcal{B}_{\psi}(f,f^{\prime}_{t-1})\right\},$ then, $\forall u\in\Omega$,
$\langle f^{\ast}-u,x\rangle\leq\mathcal{B}_{\psi}(u,f^{\prime}_{t-1})-\mathcal{B}_{\psi}(u,f^{\ast})-\mathcal{B}_{\psi}(f^{\ast},f^{\prime}_{t-1}).$ (31)
Lemma 2 (Self-confident tuning [41])
Let $\{x_{t}\}_{t=1}^{n}$ be a sequence with $x_{t}\in[0,B]$ for all $t$. Then
$\sum_{t=1}^{n}\frac{x_{t}}{\sqrt{1+\sum_{s=1}^{t-1}x_{s}}}\leq 4\sqrt{1+\sum_{t=1}^{n}x_{t}}+B.$ (32)
Let $\psi_{t}(\mathbf{G})=\frac{1}{2\eta_{t}}\|\mathbf{G}\|^{2}_{F}$. According to Lemma 1, we can obtain
~$\langle\nabla\tilde{\ell}_{t}(\mathbf{G}_{t}),\mathbf{G}_{t}-\mathbf{G}\rangle$ (33)
$=\langle\mathbf{M}_{t},\mathbf{G}_{t}-\mathbf{G}^{\prime}_{t+1}\rangle+\langle\nabla\tilde{\ell}_{t}(\mathbf{G}_{t})-\mathbf{M}_{t},\mathbf{G}_{t}-\mathbf{G}^{\prime}_{t+1}\rangle$
+~$\quad+\langle\nabla\tilde{\ell}_{t}(\mathbf{G}_{t}),\mathbf{G}^{\prime}_{t+1}-\mathbf{G}\rangle$
$\leq\mathcal{B}_{\psi_{t}}(\mathbf{G}^{\prime}_{t+1},\mathbf{G}^{\prime}_{t})-\mathcal{B}_{\psi_{t}}(\mathbf{G}^{\prime}_{t+1},\mathbf{G}_{t})-\mathcal{B}_{\psi_{t}}(\mathbf{G}_{t},\mathbf{G}^{\prime}_{t})$
+~$\quad+\langle\nabla\tilde{\ell}_{t}(\mathbf{G}_{t})-\mathbf{M}_{t},\mathbf{G}_{t}-\mathbf{G}^{\prime}_{t+1}\rangle$
$\quad+\mathcal{B}_{\psi_{t}}(\mathbf{G},\mathbf{G}^{\prime}_{t})-\mathcal{B}_{\psi_{t}}(\mathbf{G},\mathbf{G}^{\prime}_{t+1})-\mathcal{B}_{\psi_{t}}(\mathbf{G}^{\prime}_{t+1},\mathbf{G}^{\prime}_{t}),$
$\leq\mathcal{B}_{\psi_{t}}(\mathbf{G},\mathbf{G}^{\prime}_{t})-\mathcal{B}_{\psi_{t}}(\mathbf{G},\mathbf{G}^{\prime}_{t+1})$
+~$\quad+\langle\nabla\tilde{\ell}_{t}(\mathbf{G}_{t})-\mathbf{M}_{t},\mathbf{G}_{t}-\mathbf{G}^{\prime}_{t+1}\rangle-\mathcal{B}_{\psi_{t}}(\mathbf{G}^{\prime}_{t+1},\mathbf{G}_{t})$
$=\mathcal{B}_{\psi_{t}}(\mathbf{G},\mathbf{G}^{\prime}_{t})-\mathcal{B}_{\psi_{t}}(\mathbf{G},\mathbf{G}^{\prime}_{t+1})+\frac{\eta_{t}}{2}\|\nabla\tilde{\ell}_{t}(\mathbf{G}_{t})-\mathbf{M}_{t}\|_{F}^{2}$
$\quad-\frac{1}{2\eta_{t}}\|\mathbf{G}^{\prime}_{t+1}-\mathbf{G}_{t}-\eta_{t}(\nabla\tilde{\ell}_{t}(\mathbf{G}_{t})-\mathbf{M}_{t})\|_{F}^{2}$
$\leq\mathcal{B}_{\psi_{t}}(\mathbf{G},\mathbf{G}^{\prime}_{t})-\mathcal{B}_{\psi_{t}}(\mathbf{G},\mathbf{G}^{\prime}_{t+1})+\frac{\eta_{t}}{2}\|\nabla\tilde{\ell}_{t}(\mathbf{G}_{t})-\mathbf{M}_{t}\|_{F}^{2}$
Summing over $t=1,2,...,T$, we have
$\sum_{t=1}^{T}\langle\nabla\tilde{\ell}_{t}(\mathbf{G}_{t}),\mathbf{G}_{t}-\mathbf{G}\rangle$ (34)
$\leq\underbrace{\sum_{t=1}^{T}\frac{1}{2\eta_{t}}\Big{(}\|\mathbf{G}-\mathbf{G}^{\prime}_{t}\|_{F}^{2}-\|\mathbf{G}-\mathbf{G}^{\prime}_{t+1}\|_{F}^{2}\Big{)}}_{\mbox{term (a)}}$
$\quad+\underbrace{\sum_{t=1}^{T}\frac{\eta_{t}}{2}\|\nabla\tilde{\ell}_{t}(\mathbf{G}_{t})-\mathbf{M}_{t}\|_{F}^{2}}_{\mbox{term (b)}}.$
In the following, we will bound the two terms on the right-hand side respectively. First, we analyze the term (a). Notice that $\eta_{t}\leq\eta_{t-1}$ and $\mathbf{G}_{t}\in\mathcal{G}$ satisfies $\|\mathbf{G}_{t}\|_{F}^{2}\leq D$, so we have
term (a) (35)
$\leq\sum_{t=2}^{T}\Big{(}\frac{1}{2\eta_{t}}-\frac{1}{2\eta_{t-1}}\Big{)}\|\mathbf{G}-\mathbf{G}^{\prime}_{t}\|_{F}^{2}+\frac{1}{2\eta_{1}}\|\mathbf{G}-\mathbf{G}^{\prime}_{1}\|_{F}^{2}$
$\leq\frac{D^{2}}{2\eta_{T}}=D\sqrt{1+\sum_{\tau=1}^{T}\|\nabla\tilde{\ell}_{\tau}(\mathbf{G}_{\tau})-\mathbf{M}_{\tau}\|_{F}^{2}}.$
Next, we focus on the term (b). Notice that $\mathbf{G}_{0}$ is a zero matrix, and ~$\|\nabla\tilde{\ell}_{1}(\mathbf{G}_{1})\|_{F}^{2}\leq(2(M-1)d^{*}L)^{2}$. Let $\xi=2(M-1)d^{*}L$. We can bound term (b) as
term (b) (36)
$=\sum_{t=2}^{T}\frac{\eta_{t}}{2}(\|\nabla\tilde{\ell}_{t}(\mathbf{G}_{t})-\mathbf{M}_{t}\|_{F}^{2})+\|\nabla\tilde{\ell}_{1}(\mathbf{G}_{1})\|_{F}^{2}$
$\leq\sum_{t=2}^{T}\frac{\eta_{t}}{2}\Big{(}\|\nabla\tilde{\ell}_{t}(\mathbf{G}_{t})-\mathbf{M}_{t}\|_{F}^{2}\Big{)}+\xi^{2}$
$=D\cdot\sum_{t=2}^{T}\frac{\|\nabla\tilde{\ell}_{t}(\mathbf{G}_{t})-\mathbf{M}_{t}\|_{F}^{2}}{2\sqrt{1+\sum_{\tau=2}^{t}\|\nabla\tilde{\ell}_{t}(\mathbf{G}_{t})-\mathbf{M}_{t}\|_{F}^{2}}}+\xi^{2}$
$\stackrel{{\scriptstyle\text{(a)}}}{{\leq}}2D\sqrt{1+\sum_{t=2}^{T}\|\nabla\tilde{\ell}_{t}(\mathbf{G}_{t})-\mathbf{M}_{t}\|_{F}^{2}}+(2D+1)\xi^{2},$
where (a) follows from the Lemma 2.
Then we substitute the two bounds above into Eq. (34) and we have
$\sum_{t=1}^{T}\langle\nabla\tilde{\ell}_{t}(\mathbf{G}_{t}),\mathbf{G}_{t}-\mathbf{G}\rangle$ (37)
$\leq 3D\sqrt{1+\sum_{t=2}^{T}\|\nabla\tilde{\ell}_{t}(\mathbf{G}_{t})-\mathbf{M}_{t}\|_{F}^{2}}+(2D+1)\xi^{2}.$
Now we focus on the term $\sum_{t=2}^{T}\|\nabla\tilde{\ell}_{t}(\mathbf{G}_{t})-\mathbf{M}_{t}\|_{F}^{2}$. Recall that $r_{t}=2d^{*}DL+\frac{1}{\sqrt{T}}$, thus the subgradient ~$\nabla\tilde{\ell}_{t}(\mathbf{G}_{t})$ defined in Eq. (16) becomes
~$\nabla\tilde{\ell}_{t}(\mathbf{G}_{t})$ =$\stackrel{{\scriptstyle\text{(a)}}}{{=}}\frac{2}{M}\sum_{j=1}^{M}\sum^{M}_{j^{\prime}\neq j}(\mathbf{x}^{j^{\prime}}-\mathbf{x}^{j})(\mathbf{y}_{t}^{j})^{T},$ (38)
where (a) holds since $r_{t}-(\|\mathbf{x}^{j^{\prime}}-\mathbf{G}_{t}\mathbf{y}_{t}^{j}\|^{2}_{2}-\|\mathbf{x}^{j}-\mathbf{G}_{t}\mathbf{y}_{t}^{j}\|^{2}_{2})=r_{t}-2\langle\mathbf{x}^{j}-\mathbf{x}^{j^{\prime}},\mathbf{G}_{t}\mathbf{y}_{t}^{j}\rangle>0$. The corresponding reason is presented as follows
$r_{t}-2\langle\mathbf{x}^{j}-\mathbf{x}^{j^{\prime}},\mathbf{G}_{t}\mathbf{y}_{t}^{j}\rangle$ $\stackrel{{\scriptstyle\text{(a)}}}{{\geq}}r_{t}-2\|\mathbf{x}^{j}-\mathbf{x}^{j^{\prime}}\|_{2}\|\mathbf{G}_{t}\mathbf{y}_{t}^{j}\|_{2}$ (39)
$\stackrel{{\scriptstyle\text{(b)}}}{{\geq}}r_{t}-2\|\mathbf{x}^{j}-\mathbf{x}^{j^{\prime}}\|_{2}\|\mathbf{G}_{t}\|_{2}\|\mathbf{y}_{t}^{j}\|_{2}$
$\stackrel{{\scriptstyle\text{(c)}}}{{\geq}}r_{t}-2\|\mathbf{x}^{j}-\mathbf{x}^{j^{\prime}}\|_{2}\|\mathbf{G}_{t}\|_{F}\|\mathbf{y}_{t}^{j}\|_{2}$
$\geq r_{t}-2d^{*}DL>0,$
where (a) holds based on the Cauchy-Schwarz inequality, (b) follows from the definition of spectral norm, and (c) follows from the fact that $\|\mathbf{G}_{t}\|_{2}\leq\|\mathbf{G}_{t}\|_{F}$.
Hence, $\mathbbm{1}^{j,j^{\prime}}_{t}(r_{t})$ always equals to $1$ and we have
~$\nabla\tilde{\ell}_{t}(\mathbf{G}_{t})-\mathbf{M}_{t}$ =~$=\nabla\tilde{\ell}_{t}(\mathbf{G}_{t})-\nabla\tilde{\ell}_{t-1}(\mathbf{G}_{t-1})$ (40)
$=\frac{2}{M}\sum_{j=1}^{M}\sum^{M}_{j^{\prime}\neq j}(\mathbf{x}^{j^{\prime}}-\mathbf{x}^{j})(\mathbf{y}_{t}^{j}-\mathbf{y}_{t-1}^{j})^{T}.$
Then we have
$\sum_{t=1}^{T}\langle\nabla\tilde{\ell}_{t}(\mathbf{G}_{t}),\mathbf{G}_{t}-\mathbf{G}\rangle$ (41)
$\leq 3D+(2D+1)\xi^{2}+6Dd^{*}\sqrt{M\sum_{t=2}^{T}\sum_{j\in[M]}\big{\|}\mathbf{y}_{t}^{j}-\mathbf{y}_{t-1}^{j}\big{\|}_{2}^{2}}$
Based on this result, we take the expectation of Eq. (41) and apply Jensen’s inequality to have
$\sum_{t=1}^{T}\mathbb{E}\big{[}\langle\nabla\tilde{\ell}_{t}(\mathbf{G}_{t}),\mathbf{G}_{t}-\mathbf{G}\rangle\big{]}$ (42)
$\leq 3D+(2D+1)\xi^{2}+6Dd^{*}\sqrt{M\sum_{t=2}^{T}\sum_{j\in[M]}\mathbb{E}\big{\|}\mathbf{y}_{t}^{j}-\mathbf{y}_{t-1}^{j}\big{\|}_{2}^{2}}$
$=\mathcal{O}\Big{(}\sqrt{M\sum_{t\in[T]}\sum_{j\in[M]}\mathbb{E}\big{\|}\mathbf{y}_{t}^{j}-\mathbf{y}_{t-1}^{j}\big{\|}_{2}^{2}}\Big{)}.$
Based on the theoretical findings in Eq. (14), we have
$\frac{1}{T}\sum_{t\in[T]}\mathbb{P}_{t}(\mathbf{G}_{t})$ (43)
$\leq\frac{1}{T}\sum_{t\in[T]}\mathbb{E}[\tilde{\ell}_{t}(\mathbf{G}_{t})]$
$\stackrel{{\scriptstyle\text{(a)}}}{{\leq}}\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}[\langle\nabla\tilde{\ell}_{t}(\mathbf{G}_{t}),\mathbf{G}_{t}-\mathbf{G}\rangle]+\frac{1}{T}\sum_{t\in[T]}\mathbb{E}[\tilde{\ell}_{t}(\mathbf{G})]$
$\leq\mathcal{O}\Big{(}\frac{1}{T}\sqrt{M\sum_{t\in[T]}\sum_{j\in[M]}\mathbb{E}\big{\|}\mathbf{y}_{t}^{j}-\mathbf{y}_{t-1}^{j}\big{\|}_{2}^{2}}\Big{)}+\frac{1}{T}\sum_{t\in[T]}\mathbb{E}[\tilde{\ell}_{t}(\mathbf{G})],$
where (a) holds based on the fact that for any $\mathbf{G}\in\mathcal{G}$, ~$\tilde{\ell}_{t}(\mathbf{G}_{t})-\tilde{\ell}_{t}(G)\leq\langle\nabla\tilde{\ell}_{t}(\mathbf{G}_{t}),\mathbf{G}_{t}-\mathbf{G}\rangle$. This fact holds since the surrogate loss ~$\tilde{\ell}_{t}(\mathbf{G}_{t})$ is a convex function w.r.t. $\mathbf{G}_{t}$.
This completes the proof of Theorem 1
-B Proof of Corollary 1
Based on the fact that $\mathbf{y}_{t}^{j}=\mathbf{H}_{t}\mathbf{x}^{j}+W_{t}$ and $\|\mathbf{x}^{j}\|_{2}\leq\gamma_{X}^{2}$, we find that $\mathbb{E}\|\mathbf{y}_{t}^{j}-\mathbf{y}_{t-1}^{j}\|_{2}^{2}$ satisfies
$\mathbb{E}\|\mathbf{y}_{t}^{j}-\mathbf{y}_{t-1}^{j}\|_{2}^{2}$ (44)
=$\stackrel{{\scriptstyle\text{(a)}}}{{=}}\mathbb{E}\|(\mathbf{H}_{t}-\mathbf{H}_{t-1})\mathbf{x}^{j}\|_{2}^{2}+\mathbb{E}\|W_{t}-W_{t-1}\|_{2}^{2}$
$\leq\gamma^{2}_{X}\mathbb{E}\|\mathbf{H}_{t}-\mathbf{H}_{t-1}\|_{F}^{2}+\mathbb{E}\|W_{t}-W_{t-1}\|_{2}^{2}$
$=\gamma^{2}_{X}\mathbb{E}\|\mathbf{H}_{t}-\mathbf{U}_{t}+\mathbf{U}_{t}-\mathbf{U}_{t-1}+\mathbf{H}_{t-1}-\mathbf{U}_{t-1}\|_{F}^{2}$
$\quad+\mathbb{E}\|W_{t}-\mathbb{E}[W]+\mathbb{E}[W]-W_{t-1}\|_{2}^{2}$
$\leq 3\gamma^{2}_{X}\Big{(}\mathbb{E}\|\mathbf{H}_{t}-\mathbf{U}_{t}\|_{F}^{2}+\mathbb{E}\|\mathbf{H}_{t-1}-\mathbf{U}_{t-1}\|_{F}^{2}$
$\quad+\mathbb{E}\|\mathbf{U}_{t}-\mathbf{U}_{t-1}\|_{F}^{2}\Big{)}+2\sigma_{W}^{2}.$
where (a) follows from the fact that the additive white Gaussian noise is zero-mean, and is independent of the channel gain and input.
Based on the proof of Theorem 1 and the physical quantities introduced in Section IV, we have
$\sum_{t=1}^{T}\mathbb{E}[\langle\nabla\tilde{\ell}_{t}(\mathbf{G}_{t}),\mathbf{G}_{t}-\mathbf{G}\rangle]$ (45)
$\leq 3D+(2D+1)\xi^{2}+6Dd^{*}(M-1)\Big{(}\gamma_{X}\sqrt{6(\sigma^{\mathbf{H}}_{1:T})^{2}}$
$\quad+\gamma_{X}\sqrt{3(\Sigma^{\mathbf{H}}_{1:T})^{2}}+\sqrt{2\sigma_{W}^{2}T}\Big{)}$
$=\mathcal{O}\Big{(}M\Big{(}\sqrt{(\sigma^{\mathbf{H}}_{1:T})^{2}}+\sqrt{(\Sigma^{\mathbf{H}}_{1:T})^{2}}+\sqrt{\sigma_{W}^{2}T}\Big{)}\Big{)}.$
This completes the proof of Corollary 1.
-C Proof of Theorem 2
The expected error probability $\mathbb{P}(\bar{\mathbf{G}})$ of the final output $\bar{\mathbf{G}}=\frac{1}{T}\sum_{t\in[T]}\mathbf{G}_{t}$ of Algorithm 2 satisfies
$\mathbb{P}(\bar{\mathbf{G}})$ $\leq\frac{1}{M}\sum_{j=1}^{M}\sum^{M}_{j^{\prime}\neq j}P\Big{\{}\|\mathbf{x}^{j^{\prime}}-\bar{\mathbf{G}}\mathbf{y}^{j}\|^{2}_{2}\leq\|\mathbf{x}^{j}-\bar{\mathbf{G}}\mathbf{y}^{j}\|^{2}_{2}\Big{\}}$ (46)
$\leq\frac{1}{M}\sum_{j=1}^{M}\sum^{M}_{j^{\prime}\neq j}\mathbb{E}\big{[}r-\|\mathbf{x}^{j^{\prime}}-\bar{\mathbf{G}}\mathbf{y}^{j}\|^{2}_{2}+\|\mathbf{x}^{j}-\bar{\mathbf{G}}\mathbf{y}^{j}\|^{2}_{2}\big{]}_{+},$
where $r=2Dd^{*}L+\frac{1}{\sqrt{T}}\geq 1$ since we choose $d^{*}\geq\frac{1}{2DL}$.
Then we denote $[r-\|\mathbf{x}^{j^{\prime}}-\bar{\mathbf{G}}\mathbf{y}^{j}\|^{2}_{2}+\|\mathbf{x}^{j}-\bar{\mathbf{G}}\mathbf{y}^{j}\|^{2}_{2}]_{+}$ as ~$\tilde{\ell}^{j,j^{\prime}}(\bar{\mathbf{G}})$ and we have
$\frac{1}{M}\sum_{j=1}^{M}\sum^{M}_{j^{\prime}\neq j}\mathbb{E}[\tilde{\ell}^{j,j^{\prime}}(\bar{\mathbf{G}})]$ $=\frac{1}{M}\sum_{j=1}^{M}\sum^{M}_{j^{\prime}\neq j}\mathbb{E}[\tilde{\ell}^{j,j^{\prime}}(\frac{1}{T}\sum_{t\in[T]}\mathbf{G}_{t})]$ (47)
$\stackrel{{\scriptstyle\text{(a)}}}{{\leq}}\frac{1}{MT}\sum^{M}_{j=1}\sum^{M}_{j^{\prime}\neq j}\sum_{t\in[T]}\mathbb{E}[\tilde{\ell}^{j,j^{\prime}}(\mathbf{G}_{t})]$
$=\frac{1}{MT}\sum^{M}_{j=1}\sum^{M}_{j^{\prime}\neq j}\sum_{t\in[T]}\mathbb{E}\Big{[}\tilde{\ell}^{j,j^{\prime}}(\mathbf{G}_{t})$
~$\quad-\tilde{\ell}_{t}^{j,j^{\prime}}(\mathbf{G}^{*})+\tilde{\ell}_{t}^{j,j^{\prime}}(\mathbf{G}^{*})\Big{]}$
$=\frac{\mathbb{E}[\mbox{Reg}_{T}]}{T}+\frac{1}{T}\sum_{t\in[T]}\mathbb{E}[\tilde{\ell}_{t}(\mathbf{G}^{*})],$
where (a) holds based on Jensen’s inequality and the fact that ~$\tilde{\ell}^{j,j^{\prime}}(\bar{\mathbf{G}})$ is a convex function w.r.t $\bar{\mathbf{G}}$.
Thus we have
$\mathbb{P}(\bar{\mathbf{G}})-\frac{1}{T}\sum_{t\in T}\mathbb{E}[\tilde{\ell}_{t}(\mathbf{G}^{*})]$ $=\mathcal{O}\Big{(}M\sqrt{\frac{\sigma_{\mathbf{H}}^{2}}{T}}+M\sqrt{\frac{\sigma_{W}^{2}}{T}}\Big{)}.$ (48)
This completes the proof of Theorem 2.
-D Proof of Theorem 3
Lemma 3 (Corollary 9 in [23])
For the optimistic OMD with the log-barrier regularizer, if the hint $\mathbf{m}_{t}$ satisfies $\mathbf{m}_{t,i}=\boldsymbol{\ell}_{\alpha_{i}(t),i}$, the loss estimator $\hat{\boldsymbol{\ell}}$ satisfies ^$\hat{\boldsymbol{\ell}}_{t,i}=\frac{(\boldsymbol{\ell}_{t,i}-\mathbf{m}_{t,i})\mathbbm{1}\{i_{t}=i\}}{\mathbf{w}_{t,i}}+\mathbf{m}_{t,i}$, and the learning rate $\eta_{t}$ satisfies $\eta_{t,i}=\eta\leq\frac{1}{162}$, using the doubling trick, and we can achieve
$\mathbb{E}[\mbox{Reg}_{T}]=\tilde{\mathcal{O}}\Big{(}\sqrt{\sum_{i\in[N]}\sum_{t\in[T]}|\boldsymbol{\ell}_{t,i}-\boldsymbol{\ell}_{t-1,i}|}\Big{)},$ (49)
where we take the expectation on the regret over the randomness of algorithm.
Based on Lemma 3, taking the expectation on Eq. (49) over the channel noise and applying Jensen’s inequality lead to
$\mathbb{E}[\mbox{Reg}_{T}]\leq\tilde{\mathcal{O}}\Big{(}\sqrt{\sum_{i\in[N]}\sum_{t\in[T]}\mathbb{E}|\boldsymbol{\ell}_{t,i}-\boldsymbol{\ell}_{t-1,i}|}\Big{)}.$ (50)
For the term $\mathbb{E}|\boldsymbol{\ell}_{t,i}-\boldsymbol{\ell}_{t-1,i}|$, we can have
$\mathbb{E}|\boldsymbol{\ell}_{t,i}-\boldsymbol{\ell}_{t-1,i}|$ (51)
$=\mathbb{E}\Big{|}\frac{1}{M}\sum_{j\in[M]}\big{(}\mathbbm{1}_{t}^{j}(i)-\mathbbm{1}_{t-1}^{j}(i)\big{)}\Big{|}$
$\leq\frac{1}{M}\sum_{j\in[M]}\mathbb{E}\Big{|}\mathbbm{1}_{t}^{j}(i)-\mathbbm{1}_{t-1}^{j}(i)\Big{|}$
$\leq\frac{1}{M}\sum_{j\in[M]}\big{[}\mathbb{E}\big{|}\mathbbm{1}_{t}^{j}(i)-\mathbb{E}[\mathbbm{1}_{t}^{j}(i)]\big{|}+\mathbb{E}\big{|}\mathbbm{1}_{t-1}^{j}(i)$
$\quad-\mathbb{E}[\mathbbm{1}_{t-1}^{j}(i)]\big{|}+\big{|}\mathbb{E}[\mathbbm{1}_{t}^{j}(i)]-\mathbb{E}[\mathbbm{1}_{t-1}^{j}(i)]\big{|}\big{]}.$
We then upper-bound the term $\mathbb{E}|\mathbbm{1}_{t}^{j}(i)-\mathbb{E}[\mathbbm{1}_{t}^{j}(i)]|$ as
$\mathbb{E}\big{|}\mathbbm{1}_{t}^{j}(i)-\mathbb{E}[\mathbbm{1}_{t}^{j}(i)]\big{|}$ $=\mathbb{E}\sqrt{\big{(}\mathbbm{1}_{t}^{j}(i)-\mathbb{E}[\mathbbm{1}_{t}^{j}(i)]\big{)}^{2}}$ (52)
$\stackrel{{\scriptstyle\text{(a)}}}{{\leq}}\sqrt{\mathbb{E}\big{(}\mathbbm{1}_{t}^{j}(i)-\mathbb{E}[\mathbbm{1}_{t}^{j}(i)]\big{)}^{2}}=\sigma[\mathbbm{1}_{t}^{j}(i)],$
where (a) holds based on Jensen’s inequality and $\sigma[\mathbbm{1}_{t}^{j}(i)]$ denotes the standard deviation of random variable $\mathbbm{1}_{t}^{j}(i)$.
To sum up, we have
$\frac{1}{T}\sum_{t\in[T]}\mathbb{P}(C_{t})$ (53)
$\leq\frac{1}{T}\mathbb{E}[\mbox{Reg}_{T}]+\frac{1}{T}\sum_{t\in[T]}\ell_{t}(C^{*})$
$\leq\tilde{\mathcal{O}}\Big{(}\frac{1}{T}\sqrt{\sum_{i\in[N]}\sum_{t\in[T]}\frac{1}{M}\sum_{j\in[M]}\big{|}\mathbb{P}^{j}_{t}(i)-\mathbb{P}^{j}_{t-1}(i)\big{|}}$
$+\frac{1}{T}\sqrt{\sum_{i\in[N]}\sum_{t\in[T]}\sum_{j\in[M]}\frac{\sigma[\mathbbm{1}_{t}^{j}(i)]}{M}}\Big{)}+\frac{1}{T}\sum_{t\in[T]}\ell_{t}(C^{*}),$
where $\mathbb{P}_{t}^{j}(i)=\mathbb{E}[\mathbbm{1}_{t}^{j}(i)]$.
This completes the proof of Theorem 3.
-E Proof of Corollary 2
According to the proof of Theorem 3, we can directly derive the upper bound of the regret at this scenario where the number $M$ of codewords satisfies $M=2$ as follows.
$\mathbb{E}[\mbox{Reg}_{T}]$ $=\tilde{\mathcal{O}}\Big{(}\sqrt{\sum_{i\in[N]}\sum_{t\in[T]}\sum_{j=1}^{2}\frac{\sigma[\mathbbm{1}_{t}^{j}(i)]}{2}}$ (54)
$\quad+\sqrt{\sum_{i\in[N]}\sum_{t\in[T]}\sum_{j=1}^{2}\frac{\big{|}\mathbb{P}_{t}^{j}(i)-\mathbb{P}_{t-1}^{j}(i)\big{|}}{2}}\Big{)}.$
Given that $Z_{t}$ is a zero-mean Gaussian noise with the variance of $\sigma^{2}_{Z_{t}}$ at this scenario, we can calculate the corresponding expected error probability $\mathbb{P}_{t}^{j}(i)=\mathbb{E}[\mathbbm{1}^{j}_{t}(i)]$ as
[…]
Hence, the term $\big{|}\mathbb{P}_{t}^{j}(i)-\mathbb{P}_{t-1}^{j}(i)\big{|}$ becomes
[…]
$\stackrel{{\scriptstyle\text{(a)}}}{{\leq}}\frac{1}{\sqrt{2\pi}}\Big{|}\frac{d_{i}}{2\sigma_{Z_{t}}}-\frac{d_{i}}{2\sigma_{Z_{t-1}}}\Big{|}$
$=\frac{d_{i}|\sigma_{Z_{t}}-\sigma_{Z_{t-1}}|}{2\sqrt{2\pi}\sigma_{Z_{t}}\sigma_{Z_{t-1}}},$
where (a) follows from the mean value theorem of integrals and the fact that $|\exp(-x)|\leq 1,\forall x\geq 0$.
We then focus on the standard deviation $\sigma[\mathbbm{1}^{j}_{t}(i)]$ of $\mathbbm{1}^{j}_{t}(i)$
$\sigma[\mathbbm{1}^{j}_{t}(i)]$ $=\sqrt{\mathbb{E}[\mathbbm{1}^{j}_{t}(i)]^{2}-(\mathbb{P}_{t}^{j}(i))^{2}}$ (57)
$=\sqrt{\mathbb{P}_{t}^{j}(i)\Big{(}1-\mathbb{P}_{t}^{j}(i)\Big{)}}$
$=\sqrt{Q\big{(}\frac{d_{i}}{2\sigma_{Z_{t}}}\big{)}\Big{(}1-Q\big{(}\frac{d_{i}}{2\sigma_{Z_{t}}}\big{)}\Big{)}}.$
Based on the symmetry of codewords in $2$-ary codebook, we can have
$\mathbb{E}[\mbox{Reg}_{T}]$ $=\tilde{\mathcal{O}}\Big{(}\sqrt{\sum_{i\in[N]}\sum_{t\in[T]}\sqrt{Q\big{(}\frac{d_{i}}{2\sigma_{Z_{t}}}\big{)}\Big{(}1-Q\big{(}\frac{d_{i}}{2\sigma_{Z_{t}}}\big{)}\Big{)}}}$ (58)
$\quad+\sqrt{\sum_{i\in[N]}\sum_{t\in[T]}|\sigma_{Z_{t}}-\sigma_{Z_{t-1}}|}\Big{)}.$
This completes the proof of Corollary 2.